# MMRel: A Relation Understanding Dataset and Benchmark in the MLLM Era

### v1

@@ Abstract
Despite the recent advancements in Multi-modal Large Language Models (MLLMs), understanding inter-object relations, i.e., interactions or associations between distinct objects, remains a major challenge for such models. This issue significantly hinders their advanced reasoning capabilities and is primarily due to the lack of large-scale, high-quality, and diverse multi-modal data essential for training and evaluating MLLMs. In this paper, we provide a taxonomy of inter-object relations and introduce Multi-Modal Relation Understanding (MMRel), a comprehensive dataset designed to bridge this gap by providing large-scale, high-quality and diverse data for studying inter-object relations with MLLMs. MMRel features three distinctive attributes: (i) It includes over 15K question-answer pairs, which are sourced from three distinct domains, ensuring large scale and high diversity; (ii) It contains a subset featuring highly unusual relations, on which MLLMs often fail due to hallucinations, thus are very challenging; (iii) It provides manually verified high-quality labels for inter-object relations. Thanks to these features, MMRel is ideal for evaluating MLLMs on relation understanding, as well as being used to fine-tune MLLMs to enhance relation understanding and even benefit overall performance in various vision-language tasks. Extensive experiments on various popular MLLMs validate the effectiveness of MMRel. Both MMRel dataset and the complete labeling scripts have been made publicly available.
@@ 1 Introduction
Multi-Modal Large Language Models (MLLMs) [1, 2, 3, 4, 5, 6, 7] have demonstrated impressive capabilities in various vision-language tasks [8, 9, 10, 11]. One key step in these tasks is to comprehend inter-object relations, which are defined by the interactions or associations between distinct objects in images [12]. Understanding these relations is crucial for MLLMs to perform advanced reasoning, provide accurate responses in visual question answering [8, 9], and generate detailed image captions [10, 11]. However, recent studies [13, 14, 15, 16] have shown that most existing MLLMs struggle with handling inter-object relations, as illustrated in Fig. 1. The primary reasons for this challenge include the lack of a formal definition of inter-object relations and the absence of benchmarks that feature large-scale, diverse, and high-quality vision-language data.
To fill up the gap by building a comprehensive dataset for MLLMs’ relation understanding, we first partition inter-object relations into three distinct categories, namely, spatial relations (e.g., dog is left to cat), action relations (e.g., boy eats pizza), and comparative relations (e.g., apple is smaller than watermelon). Furthermore, we define the term relation hallucinations by scenarios where the MLLM-generated inter-object relations align with common sense but deviate from the factual contents in images. For example, MLLMs responds that “a man drives a car” while the ground truth is “a man pushes a car”. In addition, we examine pioneer benchmarks that were not intended for but did allow assessing MLLMs’ relation understanding capabilities [17, 18, 19, 20, 21, 22, 23, 24, 25]. As Tab. 1 shows, most existing benchmarks are typically small-scale with homogeneous data, capture one or two categories of inter-object relations only, and lack diversity with images from a single domain.
@@
Dataset Scale Taxonomy Domain MLLM+++ Adv- Evaluation
@@
Spatial Action Comparative Real SD Dall-E Human subset
MMHAL [18] 12 $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× LLM
RAH [21] 500 $\checkmark$ $\checkmark$ $\times$× $\checkmark$ $\times$× $\times$× $\times$× $\times$× Yes/No
MERLIM [22] - $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\times$× $\times$× Yes/No
FAITHSCORE [23] - - - - $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× LLM
M-HalDetect [20] - - - - $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× Human
AMBER [17] 1.7K $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× Yes/No
SPEC [19] 3.5K $\checkmark$ $\times$× $\checkmark$ $\times$× $\checkmark$ $\times$× $\times$× $\times$× Choice
Hallusion [25] 150 $\times$× $\times$× $\checkmark$ $\checkmark$ $\times$× $\times$× $\times$× $\times$× Yes/No
MME [19] 60 $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\times$× $\times$× Yes/No
MMRel 15K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Yes/No
Figure 1: MLLMs often fail to understand inter-object relations.
With the definition of inter-object relations and relation hallucinations, we create Multi-Modal Relation Understanding (MMRel), a large-scale benchmark that comprises multi-modal data of three categories of relations (i.e. spatial, action, and comparative) that are sourced from three distinctive domains (i.e., real images, synthetic images from SDXL [26], and synthetic images from Dall-E [27]). We present sample images from MMRel in Fig. 2. To ensure the quality of the collected data, we adopt a semi-automatic pipeline that leverages MLLMs [27, 28] to generate images and annotations based on textual prompts and then verifies and corrects the generated images and annotations by human reviewers. As MLLMs are known to have the issue of hallucination (especially on relation understanding tasks), to evaluate and prevent hallucinations, we create a challenging adversarial subset in MMRel which utilizes relations that deviate from common sense to assess the relation understanding capabilities of MLLMs rigorously. All these lead to a benchmark that comprises over 15,000 question-answer pairs, each of which comes with clearly defined annotations and relation categories. To make the evaluation straightforward and easy to understand, we adopt a Yes/No framework, which has been widely used in previous datasets [17, 21, 22, 24, 25, 29].
Figure 2: Illustration of MMRel samples. MMRel covers three categories of inter-object relations, namely, spatial (orange), action (blue), and comparative (green), and its images are sourced from three domains of different styles: (a) natural images, (b) synthetic images by SDXL [26], and (c) synthetic images by Dall-E [27]. It is also featured by a challenging subset that comprise images with unusual relations as in (c). More Dall-E samples are provided in the Appendix.
MMRel can serve different purposes thanks to its large-scale, diverse, and high-quality vision-language data. One typical scenario is evaluating the relation understanding capabilities of existing MLLMs. We conduct comprehensive evaluations on this aspect, revealing that both existing MLLMs [2, 5, 7] and hallucination mitigation techniques [30, 31, 32] exhibit sub-optimal performance in understanding inter-object relations. Another typical scenario is instruction tuning of MLLMs for understanding inter-object relations [1, 2]. Extensive experiments verify the effectiveness of MMRel as both evaluation set and training data, and also demonstrate that MMRel data is beneficial not only to relation understanding but also to other perceptual tasks. All these studies affirm MMRel’s values in exploring more powerful MLLMs in terms of improving advanced perception.
@@
The contributions of this work can be summarized in three aspects. First, it systematically defines inter-object relations as well as relation hallucinations under the context of MLLMs. This lays a solid foundation for future study of inter-object relations and hallucination mitigation. Second, we create MMRel, a large-scale benchmark that comprises over 15K high-quality vision-language data that are sourced from diverse domains with clearly defined inter-object relations. To the best of our knowledge, this is the first large-scale benchmark on inter-object relations that comes with diverse and high-quality vision-language data. Third, extensive experiments demonstrate the value of MMRel for both evaluating existing MLLMs on relation understanding and fine-tuning for the enhanced power of relation understanding.
2 Reviews on Existing Datasets for Relation Understanding
@@
Research on inter-object relations boasts a rich history, with Scene Graph Generation (SGG) [33, 34] serving as a foundational area of focus. SGG is specifically dedicated to identifying objects and their interrelations within images. Leveraging specialized object detection methods [35, 36] that excel in detecting both tiny and inconspicuous objects, SGG datasets such as Visual Genome (VG) [12] are distinguished by their extremely detailed and comprehensive annotations. Following the recent successes [37], Multi-Modal Large Language Models (MLLMs) have received significant attention. In response, new datasets [14, 38] have been developed to assess the capabilities of MLLMs in relation understanding.
Figure 3: Comparing MMRel with existing datasets. Existing datasets exhibit the following limitations: (a)-(b) negative choices for action relations are implausible, thus can be easily ruled out; (c) evaluation metrics are complex and subjective; (d) definitions and taxonomies of relations are ambiguous. MMRel mitigates the limitations of existing datasets by enhancing the precision and utility of inter-object relation data.
@@
VL-Checklist [13] and ARO [38] are representative datasets that evaluate the capabilities of MLLMs to understand fine-grained information, including inter-object relations. However, these datasets inherit images and labels directly from VG [12] and adopt a binary Positive/Negative evaluation framework, as illustrated in Fig. 3(a)-(b). This evaluation approach has significant limitations, as it generates negative choices randomly without human verification. Specifically, the negative choices are generated by: (i) randomly altering the relations (e.g. man riding shirt); (ii) simply reversing the order of objects (e.g. grass is eating the horse), making such negative options semantically implausible or even impossible. Given this limitation, we hypothesize that MLLMs can easily deduce the correct answers even without seeing images, as the advanced capabilities of LLM enable MLLMs to effectively dismiss negative choices that starkly deviate from common sense.
@@
Moreover, although recent MLLMs [1, 2, 3, 4, 5, 6, 7] are versatile, their proficiency in object detection is limited due to pre-training on coarse granularity image-text pairs [39, 40]. Therefore, the overly dense and occasionally incorrect annotations from existing datasets, such as VG [12], are unsuitable for assessing MLLMs. Consequently, re-labeling some existing dataset is essential for developing new benchmarks that are more suitable for the evaluation of MLLMs’ capabilities in relation understanding. This has triggered the recent initiatives such as M-HalDetect [20] and AMBER [17] that concentrate on generating new labels. Specifically, M-HalDetect [20] incorporates manually annotated relations within lengthy descriptive sentences, with evaluations performed by human assessors (refer to Fig. 3(c)). However, this metric does not straightforwardly reflect the capabilities in relation understanding and introduces potential risks of subjective comparisons. On the other hand, AMBER [17] simplifies the classification of relationships to contact or not and employs a Yes/No evaluation framework (refer to Fig. 3(d)). While this approach is straightforward, it leads to an imprecise evaluation due to the restricted definition of relations.
@@
The limitations of existing benchmarks can thus be summarized: (i) They are small-scale and lack diversity; (ii) Their annotations are of low quality, with negative choices for action relations often being meaningless; (iii) They lack clear definitions and taxonomies of inter-object relations.
3 Multi-Modal Relation Understanding (MMRel) Dataset
@@
As analyzed in Sec. 1 and Sec. 2, existing datasets for relation understanding are generally constrained by their small-scale and homogeneous nature. This motivates us to develop a new benchmark, Multi-Modal Relation Understanding (MMRel), characterized by large-scale, diverse, and high-quality data on inter-object relations. In this section, we first describe the image sources and relation taxonomies in Sec. 3.1. Subsequently, we introduce a Semi-automatic Data Collection pipeline in Sec. 3.2, designed to amass large-scale, diverse, and high-quality inter-object relation data. Finally, we present the statistics and additional details of the MMRel dataset in Sec. 3.3.
@@ 3.1 Data Source and Relation Taxonomy
Following previous works [13, 38], we source our real image subset from the Visual Genome (VG) [12] and select synthetic images from the SPEC [19], which is built based on SDXL [26]. Specifically, we use the relative spatial subset from SPEC [19] for spatial relations and the relative size subset for comparative relations. To further diversify the MMRel, we also generate synthetic images via Dall-E [27]. Given the good clarity of annotations in SPEC [19], we adapt their labels to enhance our dataset’s utility for relation understanding. Details regarding annotation and image generation across other domains are outlined in Sec. 3.2.
@@
Moreover, we employ distinct annotation strategies for each category of relations within our taxonomy. For spatial relations, we limit the categories to left, right, on/over, and under, as other descriptors such as around or next to do not accurately capture relative inter-object positions. For action relations, our annotations are based on the capabilities of Multi-Modal Large Language Models (MLLMs), as detailed in Sec. 3.2. Comparative relations are annotated manually.
3.2 Semi-Automatic Data Collection
@@
In this section, we introduce our Semi-automatic Data Collection pipeline (SemiDC), as illustrated in Fig. 4. This pipeline is capable of annotating large-scale existing images and generating a substantial amount of high-quality synthetic images. As discussed in Sec. 2, re-labeling existing images is essential since their original labels are incompatible with MLLMs. To this end, we design SemiDC to generate high-quality relation annotations via GPT-4V [28] (refer to Fig. 4(a)) for large-scale VG dataset [12]. This process is divided into three stages: (i) Pre-processing: We selectively exclude images featuring complex scenes that pose challenges for GPT-4V [28] in generating accurate annotations. Specifically, we utilize the object labels and corresponding bounding boxes from the VG dataset [12] to identify and exclude images that contain more than 10 tiny objects, each occupying less than 1/20 of the total image area; (ii) Re-labeling via GPT-4V [28]: We employ the in-context learning paradigm [4, 41] to use GPT-4V [28] to generate relation annotations. For example, we provide an image and its accurate relation annotations as a reference, together with a query image accompanied by the textual prompt “Please generate the action relationships between salient objects in this image”. GPT-4V [28] will then produce annotations for the query image based on the text prompt. More details can be found in the Appendix; (iii) Human verification: We manually assess and correct the annotations that are generated by GPT-4V [28], to ensure the quality of the collected inter-object relation data.
@@
Figure 4: Pipeline of Semi-automatic Data Collection (SemiDC). (a) Re-labeling images from Visual Genome [12] datasets. (b) Generating synthetic images based on textual prompts.
@@
As MLLMs are known to have a bias towards positive responses [29, 42], we design MMRel to maintain a balance between “Yes” and “No” answers to mitigate potential biases in evaluation metrics [29]. Specifically, negative for spatial and comparative relations can be simply derived by reversing the order of two objects. For instance, “dog is left to cat” is the inverse of “cat is left to dog”. Nevertheless, as discussed in Sec. 2, randomly generated negative choices for action relations are implausible, which can be easily ruled out even without referring to the images. To address this issue, we first adopt GPT-3.5 [43] to generate plausible negative relations that are well aligned with common sense for given objects, and then verify the generated negative relations by manual checking. For instance, a suitable negative relation for “man eats cake” might be generated as “man bakes cake”. This design is a significant improvement over prior works [13, 38], in terms of both training and evaluation.
@@
Furthermore, we introduce generative models for generating more diverse inter-object relation data. The SemiDC pipeline for generating images (refer to Fig. 2(b)) is also divided into three stages: (i) Preparing textual prompts: We prepare textual prompts describing inter-object relations; (ii) Image generation via Dall-E-3 [27]: We utilize the prepared text prompts as inputs for DALL-E-3 [27] to generate corresponding images. For instance, if we want Dall-E [27] to generate an image depicting a crocodile on the dessert, then we should employ the prompt “a photo-realistic picture depicting an object-object relation: crocodile on dessert” as input. These textual prompts define our targeted relations and the visual style of the generated images. We consider four styles, photo-realistic, watercolor, abstract, and oil painting (refer to Fig. 2(c)), to enrich our dataset; (iii) Human verification of generated images: If the generated images do not match the input relations, we will modify the relations instead of re-generating new images. The same as annotating negative choices for VG [12], we utilize GPT-3.5 [43] to generate negative action relations. More details of SemiDC for both relabeling and generation of inter-object relation data are described in the Appendix.

### v2

# Jiahao Nie1,2    Gongjie Zhang311footnotemark: 1    Wenbin An4  Yap-Peng Tan2  Alex C. Kot2  Shijian Lu2

@@ Abstract
Though Multi-modal Large Language Models (MLLMs) have recently achieved significant progress, they often face various problems while handling inter-object relations, i.e., the interaction or association among distinct objects. This constraint largely stems from insufficient training and evaluation data for relation understanding, which has greatly impeded MLLMs in various vision-language generation and reasoning tasks. We attempt to address this challenge by introducing Multi-Modal Relation Understanding (MMRel), a benchmark that features large-scale, high-quality, and diverse data on inter-object relations. MMRel features three distinctive attributes: (i) It contains over 22K question-answer pairs, spanning three distinct domains and covering three relation categories, ensuring both scale and diversity; (ii) it provides manually verified, high-quality labels to ensure exceptional annotation accuracy; (iii) it includes adversarial cases with highly unusual relations, offering a challenging setting for evaluating relation hallucination. These features make MMRel ideal for evaluating MLLMs on relation understanding, as well as for fine-tuning MLLMs to enhance relation comprehension capability. Extensive experiments verify the effectiveness of MMRel in evaluating and enhancing MLLMs’ relation understanding capabilities. The benchmark has been released publicly.
@@ 1 Introduction
Multi-Modal Large Language Models (MLLMs) [40, 39, 63, 30, 18, 22, 4, 64] have demonstrated remarkable capabilities in various vision-language tasks [43, 44, 66, 65, 31, 42, 13, 2, 61, 14, 11, 59, 47]. One essential component while handling these tasks is to understand the complicated relation among various objects in images [70] (e.g., the interaction or association between distinct objects [27]), which is crucial for advanced reasoning accurate VQA [28, 55, 3] and detailed image captioning [33, 58]. However, recent studies [70, 19, 41, 6] reveal that most existing MLLMs tend to struggle in interpreting inter-object relations, as illustrated in Fig. 1. This challenge arises primarily from lacking a large-scale, diverse, and high-quality benchmark tailored for MLLMs in the current era.
Figure 1: MLLMs often fail to understand inter-object relations.
@@
Benchmark #Q&A Taxonomy Domain MLLM+++ Adv. Dis. Gen.
@@
Spatial Action Comparative Real SDXL Dall-E Human Subset QA QA
MMHAL [52] 12 $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\checkmark$
RAH [15] 500 $\checkmark$ $\checkmark$ $\times$× $\checkmark$ $\times$× $\times$× $\times$× $\times$× $\checkmark$ $\times$×
MERLIM [54] - $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\times$× $\times$× $\checkmark$ $\times$×
FAITHSCORE [25] - - - - $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\checkmark$
M-HalDetect [23] - - - - $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\checkmark$
AMBER [56] 1.7K $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\checkmark$ $\times$×
SPEC [45] 3.5K $\checkmark$ $\times$× $\checkmark$ $\times$× $\checkmark$ $\times$× $\times$× $\times$× $\checkmark$ $\times$×
Hallusion [37] 150 $\times$× $\times$× $\checkmark$ $\checkmark$ $\times$× $\times$× $\times$× $\times$× $\checkmark$ $\times$×
MME [45] 60 $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\times$× $\times$× $\checkmark$ $\times$×
R-Bench [60] 11.7K $\checkmark$ $\checkmark$ $\times$× $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\checkmark$ $\times$×
MMRel 22.5K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Though several benchmarks on inter-object relations have been created, they were not intended for assessing MLLMs’ relation understanding capabilities [56, 52, 45, 23, 15, 54, 25, 21, 37, 60]. Specifically, most existing benchmarks suffer from obvious limitations in data scales, relation categories, and data diversity as shown in Table 1. We address this issue by creating a comprehensive benchmark on inter-object relations, aiming to gauge and enhance MLLMs’ relation understanding capability in various multimodal tasks. Specifically, we first taxonomize inter-object relations by three distinct categories, namely, spatial relations (e.g., a dog is left to a cat), action relations (e.g., a boy eats a burger), and comparative relations (e.g., an apple is smaller than a watermelon). Additionally, we introduce relation hallucination, a new concept referring to the scenario when MLLM-generated relations align with the common sense but contradict the actual contents in images (e.g., a MLLM responds by “a man drives a car” but the actual image content is “a man pushes a car”.
The created Multi-Modal Relation Understanding benchmark (MMRel) has several desirable features. First, it is large-scale comprising multi-modal data of three types of inter-object relations (i.e., spatial, action, and comparative) as well as a challenging adversarial subset that consists of abnormal relations deviating from the common sense. All these lead to around 22,500 Q&A pairs, each of which comes with clearly defined annotations and relation categories. Second, it has great diversity where the image data were collected from three distinctive sources including real images, SDXL [46] generated images with clean background, and Dall-E [7] generated multi-style images as illustrated in Fig. 2. Third, it comes with high-quality data as collected via a semi-automatic pipeline which first generates images and annotations with MLLMs [7, 1] and then reviews and corrects them by human annotators. Last but not least, it consists of both the widely adopted discriminative questions with Yes/No answers [56, 15, 54, 21, 37, 35, 60, 71] and generative open-ended questions via LLM-assisted evaluations [40, 52].
Figure 2: Sample images from MMRel. MMRel includes three categories of inter-object relations, spatial, action, and comparative. The images are sourced from three domains: (a) real images, (b) synthetic images generated by SDXL, and (c) synthetic images generated by Dall-E. Additional Dall-E samples are included in the supplementary material.
MMRel can serve different purposes thanks to its large-scale, diverse, and high-quality multimodal data. The primary usage is to gauge existing MLLMs in relation understanding. Extensive evaluations on MMRel reveal that most general MLLMs [39, 18, 4, 62, 9], grounding-specific MLLMs [12], and even hallucination mitigation methods [17, 24, 29] exhibit sub-optimal performance in understanding inter-object relations. Another typical usage is MLLM fine-tuning for enhanced relation understanding capability [40, 39]. Our experiments show that MMRel as training data can clearly improve MLLMs’ relation understanding and related perception capabilities.
@@
The contributions of this work can be summarized in three aspects. First, we systematically define inter-object relations and relation hallucinations under the context of MLLMs, laying a solid foundation for future study of relation hallucinations and their mitigation in MLLMs. Second, we contribute MMRel, a large-scale benchmark with over 22K high-quality vision-language samples sourced from diverse domains, each with clearly defined inter-object relations. To the best of our knowledge, this is the first large-scale benchmark dedicated to inter-object relations with superior diversity and quality. Third, extensive experiments demonstrate MMRel’s effectiveness for both evaluating current MLLMs and fine-tuning them to enhance relation understanding capabilities.
2 Existing Benchmarks on Relation Understanding
@@
The study of inter-object relations boasts a rich history, with Scene Graph Generation (SGG) [26, 68, 53] serving as a foundational area of focus. SGG was collected for identifying objects and their interrelations within images. Leveraging specialized object detection methods [49, 36] that excel in detecting both tiny and inconspicuous objects, SGG benchmarks such as Visual Genome (VG) [27] are distinguished by their extremely detailed and comprehensive annotations. Following their recent successes [48], Multi-Modal Large Language Models (MLLMs) have received significant attention. This has inspired several new benchmarks [19, 67] which were collected for assessing the relation understanding capabilities of MLLMs.
@@
Specifically, VL-Checklist [70] and ARO [67] are two representative benchmarks that evaluate MLLMs’ capabilities of understanding fine-grained information, including inter-object relations. The two benchmarks inherit images and labels from VG [27] and adopt a binary Positive/Negative evaluation framework, as illustrated in Fig. 3(a)-(b). However, both benchmarks have significant limitations, as they generate negative choices randomly without human verification. Specifically, they generate negative choices by: (i) randomly altering the relations (e.g., man riding shirt); (ii) simply reversing the order of objects (e.g., grass is eating the horse), making the generated negative options semantically implausible or even impossible. With such designs, MLLMs can easily deduce the correct answers even without seeing images, as the powerful LLM enables MLLMs to effectively dismiss those negative choices that starkly deviate from the common sense.
@@
Moreover, although recent MLLMs [40, 39, 63, 30, 18, 22, 4] are versatile, their proficiency in object detection is limited due to pre-training on coarse granularity image-text pairs [51, 10]. Therefore, the overly dense objects, occasionally incorrect annotations, and unclear taxonomies from existing benchmarks, such as VG [27], are unsuitable for assessing MLLM. Consequently, re-labeling some existing benchmarks is essential for developing new benchmarks that are more suitable for evaluating relation understanding in the MLLM era. This has triggered recent initiatives such as M-HalDetect [23] and AMBER [56] that concentrate on generating new labels. Specifically, M-HalDetect [23] incorporates manually annotated relations within lengthy descriptive sentences, with evaluations performed by human assessors (refer to Fig. 3(c)). However, this metric does not straightforwardly reflect the capabilities in relation understanding and introduces potential risks of subjective comparisons. On the other hand, AMBER [56] simplifies the classification of relationships to contact or not and employs a Yes/No evaluation framework (refer to Fig. 3(d)). While this approach is straightforward, it leads to an imprecise evaluation due to the restricted definition of relations.
@@
The limitations of existing benchmarks can be summarized as: (i) They are small-scale and lack diversity; (ii) Their annotations are of low quality, with negative choices for action relations often being meaningless; (iii) They lack clear definitions and taxonomies of inter-object relations.
Figure 3: Comparison of MMRel with existing benchmarks. (a)-(b) negative choices for action relations are implausible and can be easily ruled out; (c) evaluation metrics are overly complex and subjective; (d) considering only one relation “contact”, which is insufficient and ambiguous. MMRel mitigates these limitations by: (i) providing a comprehensive taxonomy of inter-object relations, (ii) carefully designing plausible negative action relations, and (iii) adopting discriminative Yes/No and generative open-ended evaluations.
3 Multi-Modal Relation Understanding Benchmark (MMRel)
@@
As analyzed in Sec. 1 and Sec. 2, existing benchmarks for relation understanding are generally constrained by their small-scale and homogeneous nature. This motivates us to develop a new benchmark, Multi-Modal Relation Understanding (MMRel), characterized by large-scale, diverse, and high-quality data on inter-object relations. In this section, we first describe the image sources and relation taxonomies in Sec. 3.1. Subsequently, we introduce a Semi-automatic Data Collection pipeline in Sec. 3.2, designed to amass large-scale, diverse, and high-quality inter-object relation data. Finally, we present the statistics and additional details of the MMRel in Sec. 3.3.
@@ 3.1 Data Source and Relation Taxonomy
Following previous works [70, 67], we source our real image subset from the Visual Genome (VG) [27] and select synthetic images from the SPEC [45], which is based on SDXL [46]. Specifically, we use the relative spatial subset for spatial relations and the relative size subset for comparative relations. To further diversify the MMRel, we also generate synthetic images via Dall-E [7]. Given the good clarity of annotations in SPEC [45], we adapt their labels to enhance our benchmark’s utility for relation understanding. Details regarding annotation and image generation across other domains are outlined in Sec. 3.2.
@@
Moreover, we employ distinct annotation strategies for each of three relation taxonomies. For spatial relations, we limit the categories to left, right, on/over, and under, as other descriptors such as around or next to do not accurately capture relative inter-object positions. For action relations, our annotations are based on the capabilities of Multi-Modal Large Language Models (MLLMs), as detailed in Sec. 3.2. Comparative relations are annotated manually. Our provided taxonomies of relations are more suitable in the MLLM era, which is further discussed in the supplementary material. We also analyze the complexity of MMRel in the supplementary material.
3.2 Semi-Automatic Pipeline
@@
In this section, we introduce our Semi-automatic Data Collection pipeline (SemiDC), as illustrated in Fig. 4. This pipeline is capable of annotating large-scale existing images and generating a substantial amount of high-quality synthetic images. As discussed in Sec. 2, re-labeling existing images is essential since their original labels are incompatible with MLLMs. To this end, we design SemiDC to generate high-quality relation annotations via GPT-4V [1] (refer to Fig. 4(a)) for large-scale VG benchmark [27]. This process is divided into three stages: (i) Pre-processing: We selectively exclude images featuring complex scenes that pose challenges for GPT-4V [1] in generating accurate annotations. Specifically, we utilize the object labels and corresponding bounding boxes from the VG benchmark [27] to identify and exclude images that contain more than 10 tiny objects, each occupying less than 1/20 of the total image area; (ii) Re-labeling via GPT-4V [1]: We employ the in-context learning paradigm [30, 69] to use GPT-4V [1] to generate relation annotations. For example, we provide an image and its accurate relation annotations as a reference, together with a query image accompanied by the textual prompt “Please generate the action relationships between salient objects in this image”. GPT-4V [1] will then produce annotations for the query image based on the text prompt. More details can be found in the supplementary material; (iii) Human verification: We manually assess and correct the annotations that are generated by GPT-4V [1], to ensure the quality of the collected inter-object relation data.
@@
Figure 4: Pipeline of Semi-automatic Data Collection (SemiDC). (a) Re-labeling real images from the Visual Genome using GPT-4V, with human verification for correctness. (b) Generating synthetic images based on textual prompts via Dall-E, followed by human verification to ensure accuracy.
Table 2: Statistics of MMRel, left: statistics of each subset, right: properties of the complete MMRel.
Domain Real SDXL Dall-E Total
Taxonomy Spat. Act. Comp. Total Spat. Comp. Total Spat. Act. Total
#Y/N QA 4.86K 4.62K 0.66K 10.14K 1.75K 0.65K 2.40K 1.20K 1.18K 2.39K 14.93K
#Gen. QA 2.43K 2.35K 0.32K 5.10K 0.87K 0.33K 1.20K 0.61K 0.62K 1.22K 7.51K
Complete MMRel
Property Obj. Spat. Act. Comp.
Number 804 5 380 9
@@
As MLLMs are known to have a bias towards positive responses [35, 38], we design MMRel to maintain a balance between “Yes” and “No” answers to mitigate potential biases in evaluation metrics [35] for binary evaluation. Specifically, negative for spatial and comparative relations can be simply derived by reversing the order of two objects. For instance, “dog is left to cat” is the inverse of “cat is left to dog”. Nevertheless, as discussed in Sec. 2, randomly generated negative choices for action relations are implausible, which can be easily ruled out even without referring to the images. To address this issue, we first adopt GPT-3.5 [8] to generate plausible negative relations that are well aligned with common sense for given objects, and then verify the generated negative relations by manual checking. For instance, a suitable negative relation for “man eats cake” might be generated as “man bakes cake”. This design significantly improves over prior works [70, 67] in terms of both training and evaluation. In addition, we provide a potential to leverage LLM to evaluate MLLMs’ capabilities of relation understanding in open-ended questions, such as “What is the action relation between a dog and a frisbee?”.
@@
Furthermore, we generate more diverse inter-object relation data via generative models. The SemiDC pipeline for generating images (refer to Fig. 2(b)) is also divided into three stages: (i) Preparing textual prompts: We prepare textual prompts describing inter-object relations; (ii) Image generation via Dall-E [7]: We utilize the prepared text prompts as inputs for Dall-E-3 [7] to generate corresponding images. For instance, if we want Dall-E [7] to generate an image depicting a crocodile on the desert, then we should employ the prompt “a photo-realistic picture depicting an object-object relation: crocodile on desert” as input. These textual prompts define our targeted relations and the visual style of the generated images. We consider four styles, photo-realistic, watercolor, abstract, and oil painting (refer to Fig. 2(c)), to enrich our benchmark; (iii) Human verification of generated images: If the generated images do not match the input relations, we will modify the relations corresponding to generated images. The same as annotating negative choices for VG [27], we utilize GPT-3.5 [8] to generate negative action relations. More details of SemiDC for both relabeling and generation of inter-object relation data are described in the supplementary material.